HETAV PATEL

Grade 11,

equipotential surfaces in some problems are spherical in shape? If potential is given by v(x,y,z)=c(x^2+y^2+z^2).show that the force is radial? hint:(force is acting along radius vector is radial)

equipotential surfaces in some problems are spherical in shape? If potential is given by v(x,y,z)=c(x^2+y^2+z^2).show that the force is radial?
hint:(force is acting along radius vector is radial)

Grade:12th pass

1 Answers

Vikas TU
7 years ago
E lctric field is given as:
Ex = -dV/dx
Ey = -dV/dy
Ez=  – dV/dz
and Force = qE
that also symbolizes.
that
F = -qdV/dx and all other  parameters.
thus dV/dx = 2cx
dV/dy = 2cy
and
dV/dz = 2zc
Fnet = Fx + Fy +Fz
ssum it all and it would be the linear combination as all variables have power .i.e degree 1.